Grade 8General Physics

A cube of side b has a charge q at each of its vertices. Determine the potential and electric field due to this charge array at the centre of the cube?

prasanjeet kumar
12 Years agoGrade 8

1 Answer

ritu mahajan
8 Years ago
The length of diagonal of the cube of each side x is √3xsquare=x√3Distane between centre of cube and each vertex,r=x√3÷2Now,potential ,V=1/4π€°×q/rsince cube has 8 charges each of value q are present there Then,V=1/4π€°×8q÷x√3/2 V=4q/√3π€°x